OpenStudy (anonymous):

MEDAL & FAN ANYBODY ? Using the equation y = –5(x + 3)2 + 5, which statement is true, if any? a. The parabola opens upwards and the directix is vertical. b. The parabola opens downward and the directix is horizontal. c. The parabola opens to the right and the directix is vertical. d. The parabola opens to the left and the directix is horizontal.

OpenStudy (anonymous):

To find opening y=a(x-h)²+k up and down x=a(y-k)²+h right and left

OpenStudy (anonymous):

If the "x" is the squared term then the parabola is a "horizontal"if the "y" is the squared term then the parabola is a "vertical" if the number in front of the squared term is positive, it goes up or right if the number in front of the squared term is negative, it goes down or left

OpenStudy (anonymous):

Is it C @Allenismyname